OpenStudy (anonymous):

for all the numbers, x#y=xy+y. if 3#2=7#k, then k=?

OpenStudy (anonymous):

x#y= xy+y , therefore 3#2=3*2 +2 = 8 7#k=7*k+k = 8k since they are equal : 8=8k , k=1 :)

OpenStudy (anonymous):

Got it ? :)

OpenStudy (anonymous):

thank u so much.... i got it :D